OpenStudy (anonymous):

Alyssa figured the area of a circle to be 75 mm². If its radius is doubled, the new area would be _____ mm². 100 150 300 350 @ajwrstlr397

OpenStudy (anonymous):

Original radius = r Original area = pi*r^2 = 75 New radius = 2r New area = pi*(2r)^2 = pi*4*r^2 The new area is 4 times as big as the old area, so the new area is 4*75 = 300, which is answer (c).

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

You welcome :)